# Soft Manipulation Surface With Reduced Actuator Density For Heterogeneous Object Manipulation

### v1

@@ I Introduction
Figure 1: Soft Manipulation Surface: A soft fabric is attached to four actuators positioned at the corners of a 0.5 x 0.5-meter aluminium frame. The system guides an apple in a circular path using an Arduino Uno microcontroller. The overhead camera view (top left) highlights the blue region of interest, with the object’s trajectory over time shown in light blue, tracked with ArUco markers
@@
The use of a soft layer allows the manipulation of highly fragile objects without damaging them. To evaluate the advantages and limitations of this approach, we constructed a hardware setup with a 0.5 x 0.5 m actuator grid, where each actuator can move up to 0.5 m in height. The system successfully manipulated objects of various shapes, sizes, and weights, including fragile items like eggs and fruit (e.g., apples). During the experiments, one of the main challenges identified was the control of object manipulation on the soft surface, due to its non-linear behavior. Despite this, the system demonstrated strong potential for applications where gentle and precise handling of heterogeneous objects is required. To overcome the nonlinear behaviour, we tested a Reinforcement Learning (RL) policy trained in simulation and zero-shot learning to manipulate a cube on the actual hardware.
@@ II Methods
In our approach, the actuators are connected using a soft layer, which eliminates the dependency on object size. The manipulation hardware consists of four linear actuators capable of vertical movement, placed at the corners of a square (0.5x0.5 meters), each with an effective range of 0.5 meters. Each actuator is powered by a DC geared motor (RS pro 834-7590, 12v, 120 rpm) that converts rotational motion into linear motion using a pulley and belt system. A single actuator with a pulley with a diameter of 5cm and a motor with a max torque of 0.29 Nm gives a force of $\sim$ 11.6 N. The four actuators are mounted vertically on a frame composed of aluminium profiles.
@@
For precise control, we use an AS500 magnetic encoder with 12-bit resolution to monitor motor rotations. An 8-channel I2C hub addresses the limitation that all four magnetic encoder sensors share the same address. An Arduino Uno controls the system by sending PWM signals to an H-bridge that drives the motors, with a PID controller managing position control to ensure smooth and accurate movements. A 0.6 x 0.6-meter, 100% polyester soft fabric, attached to the corners of the actuators, forms a flexible, hanging surface. Most object manipulation occurs on this soft surface through pulling, rolling and sliding motions driven by the interaction between the object and the fabric. The soft surface provides significant flexibility for handling objects with different textures and levels of smoothness. Depending on the properties of the objects, the surface characteristics can be adjusted for factors such as stretchability or varying levels of friction.
An overhead camera tracks the object’s position on the fabric surface at a frequency of 12 Hz. ArUco markers placed at each corner of the 0.5 x 0.5 meter manipulation region enable accurate real-world tracking within that space. For simpler tracking, we used red-colored objects; however, with more advanced object-tracking techniques, objects of any color could be used. This setup, comprising four actuators, functions as a single module.
@@ III Experiments / Results
The following subsection describe the experimental setup and results of these experiments in more detail.
@@ IV-A Pretraining in simulation
The system uses PPO to provide position values for the actuators ($a={a_{1},a_{2},a_{3},a_{4}}$), which control the actuators’ movements learned through simulation. In the simulation, the setup consists of four linear actuators, each capable of 0.5 meters of movement along the z-axis. These actuators (shown in red) are positioned at the corners of a 1x1 meter square, connected by a flexible soft layer (shown in gray) resting on top.
@@
The policy network is trained for 300K steps in the simulation. A sphere is used for policy training. In the simulation, we also evaluated the trained agent’s success rate. To do this, the 1x1 meter region of interest was divided into 100 smaller sub-regions, each measuring 0.1x0.1 meters. A random target was assigned within each subregion, and the object was placed at a random starting position on the surface. The goal was to reach the target position without the object falling off the surface. This process was repeated 10 times for each subregion, and the mean success rate was calculated. The overall results are shown in Figure 7.
@@ IV-B Demonstration and sim-to-real
We perform a preliminary test of the policy controller on the hardware by transferring the control from the simulation to the physical setup with zero-shot learning to control an object on the hardware setup. The policy operates at a 60 Hz control frequency in the simulation, while the hardware operates at a reduced control frequency of 10 Hz due to the bottleneck caused by the camera’s position capture rate. The trained policy did not perform well in the hardware experiments due to a similar gap in the system.
@@ V Discussion
This study focuses on a single module with dimensions of 0.5 x 0.5 meters and a 0.6 x 0.6 meter cloth for manipulation, though both the module and fabric size can be adjusted as needed. To cover larger distances, actuators could be spaced further apart, but this would also require increasing each actuator’s range to maintain the desired slope. Additionally, the size of the soft surface plays a role: smaller cloths would need to stretch, while larger cloths would cause objects to “hang” excessively, limiting manipulation feasibility. The soft layer material can also be replaced to modify friction, softness, or stretchability, depending on the object type. Using multiple modules allows for better control and simultaneous manipulation of multiple objects, as modules can vary in size and fabric properties to support diverse manipulation behaviors. These modules can be replicated and combined to scale the system for larger tasks.
However, using soft layers for object handling presents the challenge of achieving robust control due to the fabric’s non-linear behavior. While reinforcement learning (RL) provides more robust control compared to classical strategies, training these models is complex. The MuJoCo simulator approximates cloth and soft body behavior using composite objects, but this simulation often suffers from a reality gap [2]. Imperfect replication of fabric folds leads to a sim-to-real gap, degrading real-world performance. Real-world RL, where policies are directly trained on hardware, could address this gap but involves high costs and lengthy training times.

### v2

@@ Abstract
footnotetext: Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them.
@@ I Introduction
Figure 1: Soft Manipulation Surface: A soft fabric is attached to four actuators positioned at the corners of a $0.5\times 0.5$-meter aluminium frame. The system guides an apple in a circular path using an Arduino Uno microcontroller. The overhead camera view (top left) highlights the blue region of interest, with the object’s trajectory over time shown in light blue, tracked with ArUco markers
@@
The use of a soft layer allows the manipulation of highly fragile objects without damaging them. To evaluate the advantages and limitations of this approach, we constructed a hardware setup with a $0.5\times 0.5$-meter actuator grid, where each actuator can move up to 0.5 m in height. The system successfully manipulated objects of various shapes, sizes, and weights, including fragile items like eggs and fruit (e.g., apples). During the experiments, one of the main challenges identified was the control of object manipulation on the soft surface, due to its non-linear behavior. Despite this, the system demonstrated strong potential for applications where gentle and precise handling of heterogeneous objects is required. To overcome the nonlinear behaviour, we tested a Reinforcement Learning (RL) policy trained in simulation and zero-shot learning to manipulate a cube on the actual hardware.
@@ II Methods
In our approach, the actuators are connected using a soft layer, which eliminates the dependency on object size. The manipulation hardware consists of four linear actuators capable of vertical movement, placed at the corners of a square ($0.5\times 0.5$-meters), each with an effective range of 0.5 meters. Each actuator is powered by a DC geared motor (RS pro 834-7590, 12v, 120 rpm) that converts rotational motion into linear motion using a pulley and belt system. A single actuator with a pulley with a diameter of 5cm and a motor with a max torque of 0.29 Nm gives a force of $\sim$ 11.6 N. The four actuators are mounted vertically on a frame composed of aluminium profiles.
@@
For precise control, we use an AS500 magnetic encoder with 12-bit resolution to monitor motor rotations. An 8-channel I2C hub addresses the limitation that all four magnetic encoder sensors share the same address. An Arduino Uno controls the system by sending PWM signals to an H-bridge that drives the motors, with a PID controller managing position control to ensure smooth and accurate movements. A $0.6\times 0.6$-meter, 100% polyester soft fabric, attached to the corners of the actuators, forms a flexible, hanging surface. Most object manipulation occurs on this soft surface through pulling, rolling and sliding motions driven by the interaction between the object and the fabric. The soft surface provides significant flexibility for handling objects with different textures and levels of smoothness. Depending on the properties of the objects, the surface characteristics can be adjusted for factors such as stretchability or varying levels of friction.
An overhead camera tracks the object’s position on the fabric surface at a frequency of 12 Hz. ArUco markers placed at each corner of the $0.5\times 0.5$-meter manipulation region enable accurate real-world tracking within that space. For simpler tracking, we used red-colored objects; however, with more advanced object-tracking techniques, objects of any color could be used. This setup, comprising four actuators, functions as a single module.
@@ III Experiments / Results
The following subsection describe the experimental setup and results of these experiments in more detail. All the experiments focus on the object’s position in 2D space (along the xy plane) and do not consider the position along the Z-axis.
@@ IV-A Pretraining in simulation
The system uses PPO to provide position values for the actuators ($a={a_{1},a_{2},a_{3},a_{4}}$), which control the actuators’ movements learned through simulation. In the simulation, the setup consists of four linear actuators, each capable of 0.5 meters of movement along the z-axis. These actuators (shown in red) are positioned at the corners of a $1\times 1$-meter square, connected by a flexible soft layer (shown in gray) resting on top.
@@
The policy network is trained for 300K steps in the simulation. A sphere is used for policy training. In the simulation, we also evaluated the trained agent’s success rate. To do this, the $1\times 1$-meter region of interest was divided into 100 smaller sub-regions, each measuring $0.1\times 0.1$-meters. A random target was assigned within each subregion, and the object was placed at a random starting position on the surface. The goal was to reach the target position without the object falling off the surface. This process was repeated 10 times for each subregion, and the mean success rate was calculated. The overall results are shown in Figure 7.
@@ IV-B Demonstration and sim-to-real
We perform a preliminary test of the policy controller on the hardware by transferring the control from the simulation to the physical setup with zero-shot learning to control an object on the hardware setup. The policy operates at a 60 Hz control frequency in the simulation, while the hardware operates at a reduced control frequency of 10 Hz due to the bottleneck caused by the camera’s position capture rate. The policies trained on spheres in the simulation were used to control the sphere, cube, and cylindrical deburrer on the hardware. We were able to move the object to the target location for some objects (Cube and Cylindrical Deburrer). However, the trained policy lacks robustness in the hardware experiments due to the sim-to-real gap and the difference in control frequencies. Figure 8 illustrates these results. Cube and Cylindrical Deburrer achieve the target position, but the actions performed are not efficient. For the sphere, the lack of control frequency resulted in the object falling out of the surface as seen in Figure 8(a). Further work is needed to increase the robustness of RL control.
(a) Sphere
(b) Cube
(c) Cylindrical Deburrer
Figure 8: Object manipulation in a physical system using a policy network. Green and red circles denote the object’s start and end positions, respectively, with blue regions as target areas. Scatter points represent observed object positions, where yellow indicates more recent observations. The line plot below shows action values based on these observations.
@@ V Discussion
This study focuses on a single module with dimensions of $0.5\times 0.5$-meters and a $0.6\times 0.6$-meter cloth for manipulation, though both the module and fabric size can be adjusted as needed. To cover larger distances, actuators could be spaced further apart, but this would also require increasing each actuator’s range to maintain the desired slope. Additionally, the size of the soft surface plays a role: smaller cloths would need to stretch, while larger cloths would cause objects to “hang” excessively, limiting manipulation feasibility. The soft layer material can also be replaced to modify friction, softness, or stretchability, depending on the object type. Using multiple modules allows for better control and simultaneous manipulation of multiple objects, as modules can vary in size and fabric properties to support diverse manipulation behaviors. These modules can be replicated and combined to scale the system for larger tasks.
However, using soft layers for object handling presents the challenge of achieving robust control due to the fabric’s non-linear behavior. While reinforcement learning (RL) provides more robust control compared to classical strategies, training these models is complex. The MuJoCo simulator approximates cloth and soft body behavior using composite objects, but this simulation often suffers from a reality gap [17]. Imperfect replication of fabric folds leads to a sim-to-real gap, degrading real-world performance. Real-world RL, where policies are directly trained on hardware, could address this gap but involves high costs and lengthy training times. Experimental findings of system limitations and dependencies of manipulation behavior on object properties on real hardware, as discussed in section III Experiments/Results, could be used to improve the RL policy’s robustness further.